OpenStudy (anonymous):

Assuming A≠0, evaluate the limit lim 3sin^2(At) / cos(At) - 1 t-->0

OpenStudy (anonymous):

Ok, we'll need to use a few identities. First, apply: \[Sin^2(\theta) = \frac{1-Cos(2 \theta)}{2}\]

OpenStudy (anonymous):

You know what, a bit shorter way: \[Sin^2(\theta) + Cos^2(\theta) = 1\] \[Sin^2(\theta) = 1-Cos^2(\theta)\] Substitute that in for your sin^2